Vectors B and D are perpendicular and have the same nonzero magnitude (B=D). Draw vector R=B-D. What is the magnitude of Vector R?
Answer #1

Magnitude of R

R^2 = B^2 + D^2

As magnitude of B = D

So

R^2 = B^2 + B^2

R = Sqrt(2) * B

==========

Comment in case any doubt, will reply for sure.. Goodluck